# Trajectory Volatility for Out-of-Distribution Detection in Mathematical Reasoning

### v1

@@ Abstract
Figure 1: Illustration of the comparisons of input/output space between mathematical reasoning and text generation scenarios. We select the MATH [7] dataset for mathematical reasoning and the OPUS [52] for text generation (translation). We use SimCSE [11] to obtain sentence embeddings and project them via UMAP [38]. Different colors represent different domains, and lighter and darker shades of the same color represent the input and output spaces, respectively. We also present input/output examples of two domains in each scenario. Appendix B.1 and B.2 show detailed settings and full examples. Projections show that input/output space overlapping of different domains under the reasoning scenario is more obvious, especially in output spaces where high-density collapse happens.
@@ 1 Introduction
Based on the availability of OOD data, detection methods can be classified into two categories. When OOD data is available, constructing a binary classifier with ID and OOD classes [9, 24, 19] is a standard method. However, the OOD data variety in real-world settings makes it impractical for classifiers to learn all OOD data features, causing scalability challenges. Thus, we focus on more generalized scenarios where OOD data is unavailable, and methods can be categorized into three main types: (1) Output-based methods assess confidence by analyzing predicted probabilities [14, 5, 34, 55, 30]; (2) Ensemble-based methods assess the uncertainty through a collection of supporting models [10, 22]; (3) Feature-based methods assess the Mahalanobis Distance between OOD samples and ID data distribution in the feature space [26, 45, 51]. Besides, there are some fragmented methods, such as gradient methods [27, 17] and auto-encoder reconstruction methods [8, 18]. However, these fragmented methods encounter optimization and computational complexity issues, leading to significant performance bottlenecks [58], so they receive relatively less attention.
Most of the above methods were proposed for classification tasks, while the research on deploying OOD detection methods on GLMs is relatively niche. Existing methods on GLMs focus on uncertainty estimation [37, 35] and embedding distance measure [46], and [46] has demonstrated that embedding-based methods are currently the only optimal solution for text generation scenarios such as summarization and translation. However, embedding-based methods encounter challenges under mathematical reasoning scenarios, as shown in Figure 1: (1) Input Space: In the text generation scenarios, input questions from different domains show more obvious clustering features, yet that under mathematical reasoning are not well distinguished. We will delve deeper into this issue in Appendix G.2; (2) Output Space: The output space shows high density under mathematical reasoning, we call this phenomenon “pattern collapse”. This is because the output space of mathematical reasoning is scalar, resulting in a compressed search space with a higher probability of overlap in two widely divergent questions. For example, “$1+3=$” and “$\int_{1}^{3}x\mathrm{d}x=$” are both “$4$”. Therefore, current methods suffer serious performance bottlenecks in mathematical reasoning scenarios.
Therefore, we transform our perspective from the static embedding representation to the dynamic embedding shift trajectory in the latent space and make the following findings: (1) Due to the pattern collapse property in the output space, trajectories corresponding to different samples are more likely to differentiate due to the hard constraints on the endpoints of trajectories compared to text generation scenarios. This theoretical perspective will be presented in Sec.2.1; (2) There exists early stabilization of GLMs for mathematical reasoning. For ID samples, GLMs have completed the main reasoning at a later stage, but this is not the case for the OOD samples. This empirical perspective will be presented in Sec.2.2. These indicate that trajectory is a promising tool to distinguish ID and OOD data.
@@
Based on these analyses, we propose the Trajectory Volatility detection algorithm (TV score) for mathematical reasoning. We conduct experiments on mainstream mathematical reasoning datasets and GLMs for the basic OOD detection, and also consider the OOD quality estimation scenario, which imposes more fine-grained requirements on the OOD scores. Results show that our method outperforms all traditional algorithms on GLMs under the mathematical reasoning scenario.
Overall, our contribution is progressively three-fold:
  • We find phenomena of GLMs under mathematical reasoning about embedding low discrepancy in the input space, pattern collapse in the output space, and early stabilization in the latent space.
  • Based on the excavated phenomena, we verify the rationality of using trajectory as OOD detection from empirical and theoretical perspectives and propose a novel detection algorithm TV Score, which distinguishes ID and OOD samples by the embedding shift trajectory volatility.
  • Our method significantly outperforms traditional algorithms on GLMs under mathematical reasoning in diverse evaluation scenarios, including offline detection, online discrimination, and quality estimation, and it can be extended to more applications like multiple-choice questions.
1.1 Problem Statement
@@
Mathematical reasoning relies on generative language models (GLMs). Let $\mathcal{X}$ be the input space and $\mathcal{Y}$ be the output space. $P_{\mathcal{X},\mathcal{Y}}$ is the joint data distribution defined over $\mathcal{X}\times\mathcal{Y}$, $P_{\mathcal{X}}$ and $P_{\mathcal{Y}}$ are the marginal distributions for inputs and outputs, respectively. GLMs are trained given input sequence ${\boldsymbol{x}}=x_{1}x_{2}...x_{t}\sim P_{\mathcal{X}}$ of length $t$ to autoregressively generate the next token in the corresponding output sequence ${\boldsymbol{y}}=y_{1}y_{2}...y_{n}\sim P_{\mathcal{Y}}$ of length $n$ over the likelihood model $p_{\boldsymbol{\theta}}({\boldsymbol{y}}|{\boldsymbol{x}})=\prod_{i=1}^{n}p_{\boldsymbol{\theta}}(y_{i}|{\boldsymbol{y}_{\prec i}},{\boldsymbol{x}})$, where each $x_{i}$ and $y_{i}$ are $d$-dimensional vectors and $\boldsymbol{\theta}$ is sampled from the parameter space $\Theta$.
@@
Assume that $\widetilde{P}_{\mathcal{X},\mathcal{Y}}$ denote a distribution sufficiently different from $P_{\mathcal{X},\mathcal{Y}}$, the goal of OOD detection in GLMs is to find a score function $f(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\theta})$ for each sample and a threshold $\epsilon$, which may rely on the features of $\mathcal{X}$, $\mathcal{Y}$, and $\Theta$, to achieve a high discrimination accuracy goal:
$\max_{f}~{}\mathbb{P}_{(\boldsymbol{x},\boldsymbol{y})\sim P_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\theta})<\epsilon\right]+\mathbb{P}_{(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}})\sim\widetilde{P}_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}},\boldsymbol{\theta})>\epsilon\right],$ (1)
and a minimum expected risk where for any two different abnormal distributions $\widetilde{P}^{\prime}_{\mathcal{X},\mathcal{Y}}$ and $\widetilde{P}^{\prime\prime}_{\mathcal{X},\mathcal{Y}}$:
$\min_{f}\left|\mathbb{P}_{(\boldsymbol{x},\boldsymbol{y})\sim\widetilde{P}^{\prime}_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\theta})>\epsilon\right]-\mathbb{P}_{(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}})\sim\widetilde{P}^{\prime\prime}_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}},\boldsymbol{\theta})>\epsilon\right]\right|.$ (2)
This is a robust goal but tough to achieve in practice, so the first goal is all current work’s main goal.
1.2 Notation and Definition
For a given sample $s_{i}$, $[{\boldsymbol{y}_{i}}]_{l}\in\mathbb{R}^{d}(0\leq l\leq L)$ denote the hidden embedding of the $l$-th layer, where $d$ is the embedding dimension and ${\boldsymbol{y}_{i}}]_{l}=\left([y_{i1}]_{l},[y_{i2}]_{l},...,[y_{id}]_{l}\right)^{\top}$. We set the embedding coordinate of the input space to $(0,[{\boldsymbol{y}_{i}}]_{0})$, the output space to (i.e., the last hidden layer) $(L,[{\boldsymbol{y}_{i}}]_{L})$, and the $L-1$ hidden layers to $\{(l,[{\boldsymbol{y}_{i}}]_{l})\}_{l=1}^{L-1}$.
Definition 1 (Component-independent Trajectory Volatility)
For sample $s_{i}$, its Component-independent Trajectory Volatility during all $L$ hidden layers is defined as the embedding difference between all neighboring layers:
${\boldsymbol{V}}(s_{i})=\frac{1}{L-1}\sum_{l=1}^{L-1}\left(|[y_{i1}]_{l+1}-[y_{i1}]_{l}|,|[y_{i2}]_{l+1}-[y_{i2}]_{l}|,...,|[y_{id}]_{l+1}-[y_{id}]_{l}|\right)^{\top}.$ (3)
Higher values of the $d$-th component of ${\boldsymbol{V}}(s_{i})$ indicate increased instability of the $d$-th dimension in the inference process for the sample $s_{i}$.
2 Motivation: Why Trajectory As The Measure?
2.1 Likelihood-High Trajectory Variation: A Macro Theoretical Perspective
@@
Figure 1 has illustrated the phenomenon of pattern collapse within the output space in mathematical reasoning scenarios, and the scalar nature of the output space makes the phenomenon intuitive. However, this special phenomenon makes detecting OOD with trajectories in mathematical reasoning scenarios more promising. We abstract this phenomenon in Figure 2, which compares the trajectory trend between different samples in the mathematical reasoning and text generation scenarios.
In mathematical reasoning, when two initial points are separated by any distance in the input space, they typically converge to a significantly closer distance in the output space after undergoing an embedding shift. However, in text generation, outputs from different samples may not exhibit this same convergence. Constraints on trajectory endpoints in mathematical reasoning allow for a greater likelihood of variation in trajectory volatility under different samples. Based on this intuition, we propose a key hypothesis:
@@
In scenarios characterized by pattern collapse in the output space, the probability of variations in the volatility of embedding shift trajectories across samples increases.
This hypothesis reflects that we are more likely to distinguish between ID and OOD samples by the trajectory volatility in mathematical reasoning than in text generation. We specify that $[{\boldsymbol{y}}_{i}]_{L}\sim\mathcal{N}(\boldsymbol{c},\Sigma^{2})$. According to the pattern collapse properties under different tasks, we can constraint the endpoint embedding $[{\boldsymbol{y}}_{i}]_{L}$ in the output space:
  • For mathematical reasoning with pattern collapse, $\Sigma\rightarrow\boldsymbol{O}$, so we approximate that $[{\boldsymbol{y}}_{i}]_{L}\equiv\boldsymbol{c}$;
@@
With such constraints, we model our key hypothesis as a main theorem:
Theorem 2.1 (Main Theorem)
We assume that $\{[{\boldsymbol{y}}_{i}]_{l}\}_{l=1}^{L}$ are all independent variables sampling from $d-$dimemsional vector space $\mathbb{R}^{d}$. For different samples $s_{i}$ and $s_{j}$, the likelihood of variations in trajectory volatility under mathematical reasoning scenarios is higher than that under text generation scenarios, which means:
$\mathbb{E}_{\{[{\boldsymbol{y}}_{i}]_{l}\}_{l=1}^{L},~{}\{[{\boldsymbol{y}}_{j}]_{l}\}_{l=1}^{L}~{}\sim~{}\mathcal{U}(\mathbb{R}^{d})}\left\{{\boldsymbol{V}}(s_{i})-{\boldsymbol{V}}(s_{j})\neq{\boldsymbol{0}}|[{\boldsymbol{y}}_{i}]_{L},[{\boldsymbol{y}}_{j}]_{L}\equiv\boldsymbol{c}\right\}$
[…]
Due to space limits, We present the complete Theoretical Proof in Appendix A. This proof can demonstrate that learning trajectories are prone to diverge between distinct samples when the output space experiences pattern collapse. This divergence increases the likelihood of discerning samples with varying features during reasoning tasks, as observed from the perspective of learning trajectories.
2.2 Early Stabilization: A Micro Empirical Perspective
Figure 3: Change curves of embedding differences between neighboring hidden layers after ID and OOD data are passed into GLMs (Llama-7B in this figure).
Figure 1 shows the minimal discrepancy in mathematical reasoning data’s input and output representations across various domains. This poses a significant risk to the efficacy of the current optimal embedding-based detection methods. However, these methods solely focus on the static nature of the embedding, neglecting its dynamic evolution within the latent space.
From this point, we first visualize the L1-norm of difference $\|\boldsymbol{y}_{l}-\boldsymbol{y}_{l-1}\|_{1}$ between the output $\boldsymbol{y}_{l}$ of each hidden layer and the output $\boldsymbol{y}_{l-1}$ of the previous layer under Llama-7B (32 layers), after ID and OOD data are input into the model, separately. Figure 3 shows the change curves within all hidden layers. We find that the embedding changes slightly until the 20th layer, and then enters the critical inference process after 20 layers. However, for ID data, the magnitude of embedding change is again suppressed after a few layers of inference, while for OOD data, the magnitude is maintained at a high level.

### v2

# Embedding Trajectory for Out-of-Distribution Detection in Mathematical Reasoning

@@ Abstract
[…]
https://github.com/Alsace08/OOD-Math-Reasoning
@@ 1 Introduction
A highly scalable detection method must not rely on specific OOD data distributions, so simulating scenarios where OOD data is unavailable is more promising. Most existing research mainly focuses on vision and text classification tasks [12, 30, 55, 26, 48]. In contrast, studies addressing OOD detection on GLMs remain relatively niche, despite the more severe risks associated with OOD perils in GLMs due to the potential for error propagation in autoregressive generated sequences [43]. Existing methods on GLMs only focus on traditional text generation scenarios like summarization and translation, and these methods did not step outside the research framework of uncertainty estimation [33, 31] and embedding distance measure [43]. Among these, [43] has demonstrated that embedding-based methods are currently the only optimal solution for text generation scenarios, they determine whether a new sample is ID or OOD by calculating the Mahalanobis Distance [22] between the new sample embedding and ID embedding distribution in the static input or output space.
Figure 1: Embedding projection and cases of input and output spaces under mathematical reasoning and text generation scenarios. We select MATH [6] dataset for mathematical reasoning and OPUS [49] for text generation, each with four diverse domains. Different colors represent different domains, with lighter and darker shades indicating input and output. We use SimCSE [9] for sentence embeddings and UMAP [34] for dimensionality reduction. Appendix B shows detailed settings and examples.
Recently, mathematical reasoning has emerged as a challenging generative task and a crucial benchmark for evaluating model abilities. However, it presents unique phenomena in both input and output spaces that render embedding-based methods inapplicable, as shown in Figure 1:
  • Input space of mathematical reasoning exhibits vague clustering features across various domains, in contrast to the more defined clusters observed in text generation. This indicates that embedding may struggle to capture the complexity of mathematical questions.
  • Output space of mathematical reasoning exhibits high-density characteristics with significant overlap between different domains. we call this phenomenon Pattern Collapse. Since the output is mathematically symbolic [53, 37], it compresses the search space, increasing the likelihood of overlap between questions from disparate domains (As cases of Figure 1a). More importantly, the sequence tokenization used in GLMs allows for substantial token sharing among mathematically distinct expressions, as these mathematical tokens are primarily drawn from digits 0-9 and finite special symbols such as decimal points and square roots, rather than diverse linguistic elements. These scalar outputs lack distinctive features associated with specific domain distributions.
More discussion of the two phenomena, especially the “pattern collapse”, will be presented in Section 6. Given the limitations of traditional methods, we aim to explore innovative OOD detection solutions in mathematical reasoning scenarios. We transform our focus from static embedding space to the dynamic embedding trajectory, this motivation stems from a theoretical insight, as shown in Figure 2: “pattern collapse” causes the convergence of the trajectory endpoints of different samples, leading to significant trajectory differences across samples. In Section 2.1, we model and prove this hypothesis to elucidate the intuition of using trajectories as a measure. Subsequently, in Section 2.2, we perform empirical experiments to investigate the underlying causes of trajectory differences between ID and OOD samples. Our findings reveal a phenomenon we term Early Stabilization, wherein GLMs achieve primary reasoning in later stages for ID samples, a pattern not observed in OOD samples. This observation provides direct evidence for the rationale behind using trajectory as a measure.
@@
Based on these analyses, in Section 3, we propose the Trajectory Volatility detection algorithm (TV Score) for mathematical reasoning. Then in Section 4 and 5, we conduct extensive OOD detection experiments with diverse datasets and GLMs to validate the effectiveness of our method. We also tackle the challenging scenario of OOD quality estimation, which raises higher precision demands on the OOD scores. Results indicate that our method surpasses all traditional algorithms for GLMs under the mathematical reasoning scenario. Additionally, we demonstrate the extension of our method to a broader range of tasks with high-density features in output spaces, such as multiple-choice questions.
Problem Statement: OOD Detection on GLMs.
@@
We start by formalizing GLMs. Let $\mathcal{X}$ and $\mathcal{Y}$ be the input and output spaces with $P_{\mathcal{X}}$ and $P_{\mathcal{Y}}$ be the marginal distributions for respective space, and $P_{\mathcal{X},\mathcal{Y}}$ is the joint data distribution defined over $\mathcal{X}\times\mathcal{Y}$. GLMs are trained given input sequence ${\bm{x}}=x_{1}x_{2}...x_{t}\sim P_{\mathcal{X}}$ of length $t$ to autoregressively generate the next token in the corresponding output sequence ${\bm{y}}=y_{1}y_{2}...y_{n}\sim P_{\mathcal{Y}}$ of length $n$ over the likelihood model $p_{\bm{\theta}}({\bm{y}}|{\bm{x}})=\prod_{i=1}^{n}p_{\bm{\theta}}(y_{i}|{\bm{y}_{\prec i}},{\bm{x}})$, where each $x_{i}$ and $y_{i}$ are taken from vocabulary $\mathcal{V}$ and $\bm{\theta}$ is sampled from the parameter space $\Theta$.
@@
Assume that $\widetilde{P}_{\mathcal{X},\mathcal{Y}}$ denote a distribution sufficiently different from $P_{\mathcal{X},\mathcal{Y}}$, the goal of OOD detection in GLMs is to find a score function $f(\bm{x},\bm{y},\bm{\theta})$ for each sample and a threshold $\epsilon$, which may rely on the features of $\mathcal{X}$, $\mathcal{Y}$, and $\Theta$, to achieve a high discrimination accuracy goal:
$\max_{f}~{}\mathbb{P}_{(\bm{x},\bm{y})\sim P_{\mathcal{X},\mathcal{Y}}}\left[f(\bm{x},\bm{y},\bm{\theta})<\epsilon\right]+\mathbb{P}_{(\bm{\widetilde{x}},\bm{\widetilde{y}})\sim\widetilde{P}_{\mathcal{X},\mathcal{Y}}}\left[f(\bm{\widetilde{x}},\bm{\widetilde{y}},\bm{\theta})>\epsilon\right].$ (1)
2 Dynamic Embedding Trajectory Distinguishes ID and OOD Samples in Mathematical Reasoning
We begin by defining the embedding trajectory and its volatility features. Given a sample $s$, we put its input to the language model $p_{\theta}$ with $L$ layers, and the output sequence consists of $T$ tokens. For the $t$-th token, we denote its output embedding at layer $l$ as $\bm{h}_{l}^{t}$, with each embedding a $d$-dimensional vector. Following the definitions in [43, 51], we define the average embedding $\bm{y}_{l}=\frac{1}{T}\sum_{t=1}^{T}\bm{h}_{l}^{t}$ as the sentence embedding at layer $l$. Then, the embedding trajectory is formed as a progressive chain of these embeddings: $\bm{y}_{0}\rightarrow\bm{y}_{1}\rightarrow\cdots\rightarrow\bm{y}_{l}\rightarrow\cdots\rightarrow\bm{y}_{L-1}\rightarrow\bm{y}_{L}$. To measure the change magnitudes of the embedding trajectory in the latent space, we define two types of trajectory volatilities:
  • Dimension-independent volatility $\bm{V}_{I}(s)\in\mathbb{R}^{d}$: For sample $s$, we first obtain the embedding difference vector $|\bm{y}_{l}-\bm{y}_{l-1}|$ between each adjacent-layer pair $l-1$ and $l$. $\bm{V}_{I}$ is the average of all differences across layers, it captures the local trajectory changes across individual dimensions:
    $\bm{V}_{I}(s)=\frac{1}{L}\cdot\sum_{l=1}^{L}|\bm{y}_{l}-\bm{y}_{l-1}|=\frac{1}{L}\cdot\sum_{l=1}^{L}\left(\left|y_{l}^{1}-y_{l-1}^{1}\right|,\cdots,\left|y_{l}^{d}-y_{l-1}^{d}\right|\right)^{\top}.$ (2)
  • Dimension-joint volatility $V_{J}(s)\in\mathbb{R}$: For sample $s$, we first obtain the L2-norm of the embedding difference $||\bm{y}_{l}-\bm{y}_{l-1}||_{2}$ between each adjacent-layer pair $l-1$ and $l$. $V_{J}$ is the average of all such differences across layers, it captures the global changes in the trajectory:
    $V_{J}(s)=\frac{1}{L}\cdot\sum_{l=1}^{L}||\bm{y}_{l}-\bm{y}_{l-1}||_{2}=\frac{1}{L}\cdot\sum_{l=1}^{L}\sqrt{\frac{1}{d}\cdot\sum_{i=1}^{d}\left(y_{l}^{i}-y_{l-1}^{i}\right)^{2}}.$ (3)
In this section, we will clarify our motivation: Why Dynamic Embedding Trajectory As The Measure? This stems from the phenomenon of “pattern collapse” in the output space, where the endpoints of different sample trajectories converge to a high-density region. This constraint potentially increases the likelihood of trajectory differences across samples. We will model and prove this theoretical intuition in Section 2.1 through the Dimension-independent volatility. After gaining insight into trajectory differences, we specifically explore the differences between ID and OOD sample trajectories. We will empirically investigate this in Section 2.2 through the Dimension-joint volatility.
2.1 Theoretical Intuition: Trajectory Differences with Higher Likelihood
@@
Figure 1 has illustrated the “pattern collapse” phenomenon in the output space in mathematical reasoning scenarios. We abstract this phenomenon in Figure 2, which compares the trajectory trend between different samples in the mathematical reasoning and text generation scenarios: In mathematical reasoning, when initial points of two trajectories are separated by any distance in the input space, they typically converge to a significantly closer distance in the output space after undergoing an embedding shift. However, in text generation, outputs from different samples may not exhibit this same convergence. This finding inspires the following theoretical intuition: Hard constraints on trajectory endpoints in mathematical reasoning allow for a higher probability of trajectory differences under different samples, as expressed by the key hypothesis:
@@
In scenarios characterized by pattern collapse in the output space, the probability of trajectory volatility differences across samples increases.
We now formalize this hypothesis. Assume that the output space is a Gaussian distribution $\mathcal{N}(\bm{c},\Sigma^{2})$, so the output embedding $\bm{y}_{L}\sim\mathcal{N}(\bm{c},\Sigma^{2})$. For two different samples $s_{i}$ and $s_{j}$, their dimension-independent volatilities are $\bm{V}_{I}(s_{i})$ and $\bm{V}_{I}(s_{j})$, respectively. We want to show that when pattern collapse exists, the probability of $\bm{V}_{I}(s_{i})-\bm{V}_{I}(s_{j})\neq\bm{0}$ will increase. According to the pattern collapse features under different scenarios, we can constraint the $\bm{y}_{L}$ as follows:
  • For mathematical reasoning with pattern collapse, $\Sigma\rightarrow\bm{O}$, so we approximate that ${\bm{y}}_{L}\equiv\bm{c}$;
  • For text generation without pattern collapse, $\Sigma=\mathrm{diag}(\delta_{1},\delta_{2},\cdots,\delta_{D})\neq\bm{O}$, so ${\bm{y}}_{L}\not\equiv\bm{c}$.
With such formalized constraints, we model our Hypothesis 1 as a main theorem:
Theorem 2.1 (Main Theorem)
We assume that $\{{\bm{y}}_{l}\}_{l=1}^{L}$ are all independent variables sampling from vector space $\mathbb{R}^{d}$. For different samples $s_{i}$ and $s_{j}$, their embedding sets are $\{[{\bm{y}}_{i}]_{l}\}_{l=1}^{L}$ and $\{[{\bm{y}}_{j}]_{l}\}_{l=1}^{L}$, respectively. The likelihood of trajectory volatility differences between $s_{i}$ and $s_{j}$ under mathematical reasoning scenarios is higher than that under text generation scenarios, which means:
$\mathbb{E}_{\{[{\bm{y}}_{i}]_{l}\}_{l=1}^{L},~{}\{[{\bm{y}}_{j}]_{l}\}_{l=1}^{L}~{}\sim~{}\mathcal{U}(\mathbb{R}^{d})}\left\{{\bm{V}}_{I}(s_{i})-{\bm{V}}_{I}(s_{j})\neq{\bm{0}}|[{\bm{y}}_{i}]_{L},[{\bm{y}}_{j}]_{L}\equiv\bm{c}\right\}$
@@
Due to space limits, we present the Complete Proof in Appendix C. This demonstrates that when pattern collapse occurs, the probability of trajectory volatility differences across samples increases.
Figure 3: Trajectory volatility curve comparisons between one ID data and ten OOD data from diverse mathematical domains. Each trajectory represents the average of all samples from the corresponding datasets, with color shading being the sample standard deviation. Llama2-7B is used for the backbone.
2.2 Empirical Analysis: Early Stabilization of ID samples
The theoretical analysis in Section 2.1 provides an intuition for using embedding trajectory as the measure for distinguishing between different samples. However, the specific trajectory differences between ID and OOD samples remain unclear. Thus, we investigate this empirically in this section.
We select eleven mathematical reasoning datasets: one for ID data and ten for OOD data. The MultiArith dataset serves as the ID data, while the OOD datasets include GSM8K, SVAMP, AddSub, SingleEq, SingleOp, and MATH. The latter encompasses five tasks across various mathematical domains: Algebra, Geometry, Counting and Probability, Number Theory, and Precalculus. This selection includes varying task types and levels of difficulty. We first train a base Llama2-7B (32 layers) using the ID training set, then inference on the ID test set and all OOD test sets. Details regarding the datasets, model, and implementation can be found in Section 4.1.
We measure the dimension-joint volatility for all samples from each ID and OOD test set. Specifically, we compute each adjacent-layer change magnitude $||\bm{y}_{l}-\bm{y}_{l-1}||_{2}$, and connect all values as a change curve, with higher value means the higher volatility. Figure 3 shows ten curve comparisons, with each sub-figure including the ID data and one OOD data. We find that the change magnitude slightly until the 20th layer. After 20 layers, for ID data, the change magnitude is again suppressed after a few layers of inference, while for OOD data, the magnitude is maintained at a relatively high level.